# Boosting Diffusion-Based Text Image Super-Resolution Model Towards Generalized Real-World Scenarios

### v2

@@ 1 Introduction
Existing SR methods for text images [25, 4, 27, 35, 26] leverage text priors to guide the restoration process, incorporating loss functions that enforce correct text structures. While these approaches perform well in many cases, they struggle with images that contain complex strokes or intricate fonts, resulting in inaccurate text feature restoration [47]. Additionally, some methods improve architectures. E.g., MACRONet [21] utilize a codebook to store character codes, which are helpful for generating structural details, and StyleGAN [16] is used to generate aesthetically pleasing text styles. However, these methods still fail when applied to LR images derived from unseen scenarios.
Recently, diffusion models (DM) [11, 32] have shown great promise in image restoration tasks, and researchers have started applying these models to text image SR [47, 31]. These methods leverage the generative capabilities of DMs to restore images by modeling both the realistic image distribution and the underlying text prior. The combination of these two modalities significantly improves image quality compared to earlier methods. However, when dealing with real-world LR images, these methods still face significant challenges, especially when fine-tuning pre-trained models. The errors inherited from the pre-training phase can hinder the model’s ability to adapt to specific downstream tasks, limiting its effectiveness. Moreover, when OCR modules generate erroneous priors for LR images, the resulting super-resolved images may contain inaccurate or nonsensical text, which degrades the quality and clarity.
Moreover, we also observe that when diffusion models are trained solely on real data, they struggle to converge effectively, preventing significant improvements in SR performance. Therefore, a dynamic sampling strategy for the real data and synthetic data (e.g., synthesized by BSRGAN [44] and Real-ESRGAN [36]) is required. These shortcomings underline the need for an SR approach that is more adaptable and capable of generalizing to various degradation levels encountered in real-world scenarios.
@@
Additionally, we argue that text image SR requires a more efficient network structure by using a SR-based diffusion rather than the large pre-trained text-to-image models like Stable Diffusion. In this paper, we employ ResShift [43] as an example to provide the SR prior. ResShift is a lightweight model that can be integrated with text modality inputs, enabling superior performance with fewer inference steps. This choice alleviates the challenges of fine-tuning large pre-trained models while maintaining high efficiency in both training and inference. We implement the interaction with text modality via the similar mechanism proposed in DiffTSR [47].
@@
By focusing on improving generalization and adaptability, our method offers a robust solution to the challenges faced in text image SR, particularly in real-world scenarios. In summary, our contributions are as follows:
@@
  • We reveal that a SR-based lightweight diffusion model is more suitable for text image SR, achieving better results and computational efficiency than the methods using large pre-trained text-to-image models. We further incorporate cross-attention from text priors and a confidence-guided mechanism, effectively enhancing text fidelity and structural consistency in super-resolution.
  • We demonstrate, through extensive experiments, that our model outperforms existing methods in the SR task for real-world text images, achieving better generalization and restoration quality.
@@ 2 Related Works
Blind Image Super-Resolution. Blind image super-resolution aims to recover high-quality images from low-resolution inputs that have undergone complex, unknown degradations in real-world scenarios. Recent advancements in this field have tackled the problem from two primary perspectives: degradation estimation [1, 9, 13, 28, 22, 33] and real-world data synthesis [3, 8, 15, 29, 20, 36, 37, 44]. The former focuses on estimating degradation models from LR images in an unsupervised manner [33], followed by the application of non-blind super-resolution methods. The latter, on the other hand, constructs LR-HR training pairs through a sophisticated degradation process that mimics real-world image deterioration. However, we observe that when it comes to text images, where characters exhibit specific and semantic text styles, relying solely on predefined degradation models is insufficient for achieving optimal restoration performance. Thus, using more advanced models like diffusion is required.
@@
Diffusion Models. Inspired by the non-equilibrium thermodynamics theory [14], diffusion model [32, 11] is proposed to model complex distributions and has achieve impressive performance in various generative modeling tasks, including image/video generation [12, 30, 38, 41, 46] and audio synthesis [19]. Diffusion model aims to build the conditional distribution through a Markov Chain, where a sequence of small Gaussian noise is gradually added to the latent vector $z$ for $T$ steps, and the sizes of noise are controlled by a variance schedule $\{\beta_{t}\}_{t=1}^{T}$. Suppose that the latent data distribution is $q(z)$, the forward process is:
$q(z_{t}|z_{t-1})=\mathcal{N}(z_{t};\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}\textbf{I}).$ (1)
By using the reparameterization tricks, $z_{t}$ at any timestep $t$ can be sampled through
$q(z_{t}|z_{0})=\mathcal{N}(z_{t};\sqrt{\bar{\alpha_{t}}z_{0},(1-\bar{\alpha_{t}}\textbf{I})}),$ (2)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha_{t}}=\Pi_{k=1}^{t}\alpha_{k}$. Since $\beta_{t}$ is small enough, the distribution of reverse $q(z_{t-1}|z_{t})$ is also Gaussian. Therefore the reverse step can also be modeled through a neural network $\theta$, denoted as $p_{\theta}(z_{t-1}|z_{t})$:
$p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t),\tilde{\beta}_{t}\textbf{I}),$ (3)
$\mu_{\theta}(z_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(z_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha_{t}}}}\epsilon_{\theta}(z_{t},t)),$ (4)
where $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$, $\epsilon_{\theta}$ is the noise predicted by neural network, and the mean $\mu_{\theta}$ is derived through Bayes rules.
@@
Benefiting from the strong power of data distribution modeling of diffusion models, plenty of works [17, 39] also achieve impressive performance in image super-resolution via the diffusion prior. However, they still have some failing cases. Text Image Super-Resolution. Text image super-resolution aims to enhance the fine details and clarity of text images, with a particular emphasis on improving text readability and recognition accuracy. Recent advances have largely focused on integrating text-specific priors and character structure priors to enhance the performance of text image SR. Specifically, prior works have leveraged text-related information from three primary perspectives: text-aware loss functions [4, 34, 35], text recognition priors [25, 40, 48], and text structure priors [21]. These approaches have demonstrated that text priors are essential for improving the structural integrity of text in the super-resolution process. Despite the effectiveness of these priors, most existing methods fail to fully exploit the available text prior information, often struggling with challenges such as diverse text styles, severe degradation, or intricate strokes. Notably, diffusion-based models have recently been applied to text image SR [31, 47]. In this paper, we propose that a SR-based diffusion prior together with a confident textural prior can achieve better performance and efficiency.
@@ 3 Methods
Figure 3: Overview of our framework. Our proposed model uses a Mixture of multi-modality module (MoM) to exchange information between the text diffusion module and the image diffusion module. Before being transmitted to the MoM module, the text feature is multiplied by the confidence score $\vec{\alpha}$ to reduce the impact of false generation. Also, $\textbf{C}_{cond_{t}}$ generated by the MoM is fused with the image feature via the Cross-attention module in the UNet module, which follows the basic layer in the pre-trained SR network.
3.1 Overview
In this paper, we propose a method to estimate the HR image $x_{0}$ from the given LR image $y$. Our method is inspired by DiffTSR [47] while has some distinct improvements. In our framework, an OCR module and a text diffusion model are also employed to recover the text information from the LR image, and a Mixture of multi-modality module (MoM) is used to facilitate the exchange of information between the image and text features.
This section offers a comprehensive description of our approach. In Section 3.2, we propose a novel data sampling strategy that progressively integrates different types of data throughout the training process. Following that, in Section 3.3, we provide an in-depth explanation of our framework, with an illustrative overview presented in Figure 3. Finally, in Section 3.4, we introduce the loss function that is used to optimize the model.
3.2 Data Sampling Strategy During Training
Our Observations. In this work, we address a key distinction from previous approaches [47, 31, 21] by focusing on improving the performance towards LR images derived from real-world scenarios. However, directly training the diffusion model with real-world LR images does not yield satisfactory results. As illustrated in Figure 2, when the SR model is trained solely on real-world LR images, the generated outputs show minimal improvement over the input LR image, indicating that the model fails to converge. Moreover, when using degraded HR images as LR inputs, as done in methods like DiffTSR [47], the results are even more problematic—producing false text and suboptimal performance. This highlights the inherent limitations of relying on the sole source of LR images during the training.
@@ 3.3 Network Architecture
Since DiffTSR [47] leverages a pre-trained text-to-image diffusion model as the backbone U-Net, it becomes difficult to eliminate the errors inherited from the pre-training phase. We think, in the context of text image SR, leveraging a pre-trained SR model is more essential for restoring the fine details in LR images, while not requiring extensive generative capabilities. A model like ResShift [43], which primarily incorporates self-attention mechanisms through the Swin Transformer architecture, offers robust spatial reasoning capabilities for enhancing image quality. However, it lacks the necessary integration of external textual information, which is crucial for tasks that involve text image restoration.
Diffusion Procedure. In this work, we aim to improve text image SR by integrating both SR priors and text-specific priors, addressing the limitations of relying solely on pre-trained text-to-image diffusion priors. While ResShift [43] is an effective SR prior, it is inherently limited in its ability to get textural priors. To overcome this, we introduce a cross-attention layer after each Swin Transformer block, enabling the model to integrate text information directly into the U-Net architecture. This cross-attention mechanism allows the model to not only perform high-quality SR but also perceive the semantic integrity of the text within the images. By doing so, the model gains the ability to restore text images with accurate and coherent textual features, addressing the unique challenges of text image restoration.
The process is carried out in the latent space, where the image $x_{0}$ is first mapped into a latent representation $z_{0}$ using VQGAN [7]. The forward process is defined as follows:
@@
where $e_{0}=y-x_{0}$ represents the residual between the LR and HR images, and $\{\eta_{t}\}_{t=1}^{T}$ is the shifting sequence that monotonically increases with timestep $t$, with $\eta_{1}\to 0$ and $\eta_{T}\to 1$. The hyper-parameter $\kappa$ controls the noise variance. The reverse distribution can be derived explicitly as [43]:
@@
Confidence Strategy. Another challenge occurs when the OCR module generates incorrect predictions for the LR image, as this may impede the image diffusion module’s ability to correct the error, leading to the generation of incorrect characters with entirely different semantic meanings. Therefore, before training the model, we utilize a TransOCR [4] to get the confidence score of the output characters. Specifically, for each character $i$ in the image, the TranOCR module outputs the prediction $\hat{i}$ and the confidence score $\alpha_{i}$. $\alpha_{i}$ could be denoted as the predicted probability of the character $i$. Then the confidence scores of the text image could be formulated as
@@
where $m$ is the length of the text. Then when we convert the text feature to MoM, we could multiply the text feature with the confidence scores of the text image, i.e.,
@@
where $z_{y}$ is the latent representation of LR image $y$ mapped by VQGAN. The method outlined above leverages the confidence scores from the TransOCR [4] module to adjust the importance of each character during training. By weighting the text features according to their confidence scores, the model can focus more on reliable characters while reducing the impact of erroneous ones.
@@ 3.4 Loss Function
First, the L1 loss is employed to minimize the pixel-wise differences between the generated super-resolved image $\tilde{x_{0}}$ and the ground truth HR image $x_{0}$. This term promotes fine-grained restoration by encouraging the model to reduce the absolute pixel differences. Then we incorporate LPIPS [45] loss which enhances the perceptual similarity between the generated and ground truth images, denoted as $\mathcal{L}_{LPIPS}$. Unlike L1 loss, which operates at the pixel level, LPIPS captures higher-level semantic differences, improving the visual quality of the output. Finally, the cross-entropy loss $\mathcal{L}_{CE}$ is introduced to ensure that the text in the generated image is semantically consistent with the ground truth text. This term is computed over the predicted characters and the true characters in the image.

### v4

@@ 1 Introduction
Existing SR methods for text images Ma et al. (2022); Chen et al. (2021); Ma et al. (2023b); Wang et al. (2020); Ma et al. (2023a) leverage text priors to guide the restoration process, incorporating loss functions that enforce correct text structures. While these approaches perform well in many cases, they struggle with images that contain complex strokes or intricate fonts, resulting in inaccurate text feature restoration Zhang et al. (2024b). Additionally, some methods improve architectures. E.g., MACRONet Li et al. (2023) utilize a codebook to store character codes, which are helpful for generating structural details, and StyleGAN Karras et al. (2020) is used to generate aesthetically pleasing text styles. However, these methods still fail when applied to LR images derived from unseen scenarios.
Recently, diffusion models (DM) Ho et al. (2020); Sohl-Dickstein et al. (2015) have shown great promise in image restoration tasks, and researchers have started applying these models to text image SR Zhang et al. (2024b); Singh et al. (2024). These methods leverage the generative capabilities of DMs to restore images by modeling both the realistic image distribution and the underlying text prior. The combination of these two modalities significantly improves image quality compared to earlier methods. However, when dealing with real-world LR images, these methods still face significant challenges. Moreover, when OCR modules generate erroneous priors for LR images, the resulting super-resolved images may contain inaccurate or nonsensical text, which degrades the quality and clarity.
Furthermore, We observe that diffusion models trained solely on real data struggle to converge effectively, limiting SR performance improvement. This necessitates a dynamic sampling strategy combining real and synthetic data (e.g., from BSRGAN Zhang et al. (2021) and Real-ESRGAN Wang et al. (2021b)). These results highlight the need for SR approaches that can better adapt to varying real-world degradation levels.
@@
Additionally, we argue that text image SR requires a more efficient network structure by using a SR-based diffusion rather than training the diffusion models from scratch. In this paper, we employ ResShift Yue et al. (2024) as an example to provide the SR prior. ResShift is a lightweight model that can be integrated with text modality inputs, enabling superior performance with fewer inference steps. This choice alleviates the challenges of fine-tuning large pre-trained models while maintaining high efficiency in both training and inference. We implement the interaction with text modality via the similar mechanism proposed in DiffTSR Zhang et al. (2024b).
@@
By focusing on improving generalization and adaptability, our method offers a robust solution to the challenges faced in text image SR. In summary, our contributions are as follows:
@@
  • We reveal that a SR-based lightweight diffusion model is more suitable for text image SR, achieving better results and computational efficiency than the methods using large latent diffusion models that is trained from scratch. We further incorporate cross-attention from text priors and a confidence-guided mechanism, effectively enhancing text fidelity and structural consistency in super-resolution.
  • Extensive experiments demonstrate that our model outperforms existing methods in the SR task for real-world text images, and achieves better generalization and restoration quality.
@@ 2 Related Works
Blind Image Super-Resolution. Blind image super-resolution aims to recover high-quality images from low-resolution inputs that have undergone complex, unknown degradations in real-world scenarios. Recent advancements in this field have tackled the problem from two primary perspectives: degradation estimation Bell-Kligler et al. (2019); Gu et al. (2019); Huang et al. (2020); Maeda (2020); Liang et al. (2022); Wang et al. (2021a) and real-world data synthesis Cai et al. (2019); Fritsche et al. (2019); Ji et al. (2020); Nazeri et al. (2019); Li et al. (2022); Wang et al. (2021b); Wei et al. (2020); Zhang et al. (2021). The former focuses on estimating degradation models from LR images in an unsupervised manner Wang et al. (2021a), followed by the application of non-blind super-resolution methods. The latter, on the other hand, constructs LR-HR training pairs through a sophisticated degradation process that mimics real-world image deterioration. However, we observe that when it comes to text images, where characters exhibit specific and semantic text styles, relying solely on predefined degradation models is insufficient for achieving optimal restoration performance. Thus, using more advanced models like diffusion is required.
@@
Diffusion Models. Inspired by the non-equilibrium thermodynamics theory Jarzynski (1997), diffusion model Sohl-Dickstein et al. (2015); Ho et al. (2020) is proposed to model complex distributions and has achieve impressive performance in various generative modeling tasks, including image/video generation Ho et al. (2022); Saharia et al. (2022); Wu et al. (2022); Ye et al. (2022); Zhang et al. (2024a) and audio synthesis Kong et al. (2020). Diffusion model aims to build the conditional distribution through a Markov Chain. Benefiting from the strong power of data distribution modeling of diffusion models, plenty of works Kawar et al. (2022); Wu et al. (2024) also achieve impressive performance in image SR via the diffusion prior. However, they still have some failing cases.
@@
Text Image Super-Resolution. Text image super-resolution aims to enhance the fine details and clarity of text images, with a particular emphasis on improving text readability and recognition accuracy. Recent advances have largely focused on integrating text-specific priors and character structure priors to enhance the performance of text image SR. Specifically, prior works have leveraged text-related information from three primary perspectives: text-aware loss functions Chen et al. (2021); Wang et al. (2019, 2020), text recognition priors Ma et al. (2022); Xu et al. (2017); Zhao et al. (2022), and text structure priors Li et al. (2023). These approaches have demonstrated that text priors are essential for improving the structural integrity of text in the super-resolution process. Despite the effectiveness of these priors, most existing methods fail to fully exploit the available text prior information, often struggling with challenges such as diverse text styles, severe degradation, or intricate strokes. Notably, diffusion-based models have recently been applied to text image SR Singh et al. (2024); Zhang et al. (2024b). In this paper, we propose that a SR-based diffusion prior together with a confident textural prior can achieve better performance and efficiency.
@@ 3 Methods
Figure 2: Overview of our framework. Our model uses the confidence score $\vec{\alpha}$ to reduce the impact of false generation. $\mathbf{C}_{cond_{t}}$ generated by MoM from Zhang et al. (2024b) is fused with image feature via the cross attention module in the Unet module, which follows the basic layer in the pre-trained SR network.
3.1 Overview
In this paper, we propose a method to estimate the HR image $x_{0}$ from the given LR image $y$. Our method is inspired by DiffTSR Zhang et al. (2024b) while has some distinct improvements. This section offers a comprehensive description of our approach. In Section 3.2, we propose a novel progressive data sampling strategy. Following that, in Section 3.3, we provide an in-depth explanation of our framework, with an illustrative overview presented in Figure 2. Finally, in Section 3.4, we introduce the loss function that is used to optimize the model.
3.2 Data Sampling Strategy During Training
Our Observations. In this work, we address a key distinction from previous approaches Zhang et al. (2024b); Singh et al. (2024); Li et al. (2023) by focusing on improving the performance towards LR images derived from real-world scenarios. However, directly training the diffusion model with real-world LR images does not yield satisfactory results. As illustrated in Figure 3, when the SR model is trained solely on real-world LR images, the generated outputs show minimal improvement over the input LR image, indicating that the model fails to converge. Moreover, when using degraded HR images as LR inputs, as done in methods like DiffTSR Zhang et al. (2024b), the results are even more problematic—producing false text and suboptimal performance. This highlights the inherent limitations of relying on the sole source of LR images during the training.
@@ 3.3 Network Architecture
We think, in the context of text image SR, leveraging a pre-trained SR model is more essential for restoring the fine details in LR images, while not requiring extensive generative capabilities. A model like ResShift Yue et al. (2024), which primarily incorporates self-attention mechanisms through the Swin Transformer architecture, offers robust spatial reasoning capabilities for enhancing image quality. However, it lacks the necessary integration of external textual information, which is crucial for tasks that involve text image restoration.
Diffusion Procedure. In this work, we aim to improve text image SR by integrating both SR priors and text-specific priors, addressing the limitations of relying solely on pre-trained text-to-image diffusion priors. While ResShift Yue et al. (2024) is an effective SR prior, it is inherently limited in its ability to get textural priors. To overcome this, we introduce a cross-attention layer after each Swin Transformer block, enabling the model to integrate text information directly into the U-Net architecture. This cross-attention mechanism allows the model to not only perform high-quality SR but also perceive the semantic integrity of the text within the images. By doing so, the model gains the ability to restore text images with accurate and coherent textual features, addressing the unique challenges of text image restoration.
The process is carried out in the latent space, where the image $x_{0}$ is first mapped into a latent representation $z_{0}$ using VQGAN Esser et al. (2021). The forward process is defined as follows:
@@
where $e_{0}=y-x_{0}$ represents the residual between the LR and HR images, and $\{\eta_{t}\}_{t=1}^{T}$ is the shifting sequence that monotonically increases with timestep $t$, with $\eta_{1}\to 0$ and $\eta_{T}\to 1$. The hyper-parameter $\kappa$ controls the noise variance. The reverse distribution can be derived explicitly as Yue et al. (2024):
@@
Confidence Strategy. Another challenge occurs when the OCR module generates incorrect predictions for the LR image, as this may impede the image diffusion module’s ability to correct the error, leading to the generation of incorrect characters with entirely different semantic meanings. Therefore, before training the model, we utilize a TransOCR Chen et al. (2021) to get the confidence score of the output characters. Specifically, for each character $i$ in the image, the TranOCR module outputs the prediction $\hat{i}$ and the confidence score $\alpha_{i}$. $\alpha_{i}$ could be denoted as the predicted probability of the character $i$. Then the confidence scores of the text image could be formulated as
@@
where $m$ is the length of the text. Then when we convert the text feature to MoM module in DiffTSR Zhang et al. (2024b), we could multiply the text feature with the confidence scores of the text image, i.e.,
@@
where $z_{y}$ is the latent representation of LR image $y$ mapped by VQGAN. The method outlined above leverages the confidence scores from the TransOCR Chen et al. (2021) module to adjust the importance of each character during training. By weighting the text features according to their confidence scores, the model can focus more on reliable characters while reducing the impact of erroneous ones.
@@ 3.4 Loss Function
First, the L1 loss is employed to minimize the pixel-wise differences between the generated super-resolved image $\tilde{x_{0}}$ and the ground truth HR image $x_{0}$. This term promotes fine-grained restoration by encouraging the model to reduce the absolute pixel differences. Then we incorporate LPIPS Zhang et al. (2018) loss which enhances the perceptual similarity between the generated and ground truth images, denoted as $\mathcal{L}_{LPIPS}$. Unlike L1 loss, which operates at the pixel level, LPIPS captures higher-level semantic differences, improving the visual quality of the output. Finally, the cross-entropy loss $\mathcal{L}_{CE}$ is introduced to ensure that the text in the generated image is semantically consistent with the ground truth text. This term is computed over the predicted characters and the true characters in the image.
@@
4 Experiments
4.1 Experimental Setup
Datasets In this work, we mainly focus on the SR of text characters in real-world scenarios as the exploration for complex characters, including Chinese characters. Following Zhang et al. (2024b), we preprocess the image datasets with the following steps: (1) remove the images with a resolution smaller than 64 pixels; (2) only retain the images with a length of texts not larger than 24; (3) only retain the images with a width-to-height greater than 2; (4) resize the image to $128\times 512$. Then, we have over 400K HR text images from the real-world scenario, and each of these images has a paired LR image. The degradation to generate $\hat{x_{0}}$ and $\hat{y}$ is proposed in BSRGAN Zhang et al. (2021). We divide the dataset into training and testing sets, with 80% of the instances divided in the training set and the remaining instances divided in the testing set. We also include the CTR-TSR-Test dataset from Yu et al. (2021) for evaluation.
Evaluation Metrics and Baselines. To assess the performance of our proposed model, we compare it with several state-of-the-art image SR methods, including general SR approaches (e.g., ESRGAN Wang et al. (2021b), NAFNet Chen et al. (2022)) and recent text image SR techniques (e.g., SRCNN Dong et al. (2015), TSRN Wang et al. (2020), TSBRN Chen et al. (2021), MACRONet Li et al. (2023), DiffTSR Zhang et al. (2024b)). When evaluating on the CTR-TSR-Test dataset, we only train our model on its corresponding training set. When testing on our collected dataset, for the baselines without training code, we adopt their released models; for these with training codes, we finetune them with our collected dataset for a fair comparison.
We utilize five evaluation metrics to quantify the effectiveness of these methods. First, we compute the peak signal-to-noise ratio (PSNR) and the learned perceptual image patch similarity (LPIPS) Zhang et al. (2018). To further assess the realism of the generated images, we report the Frechet Inception Distance (FID) Heusel et al. (2017). For a more specific evaluation of the text fidelity in restored text images, we incorporate word accuracy (ACC) and normalized edit distance (NED) Ma et al. (2023b). Notably, we use the pre-trained TransOCR Chen et al. (2021); Yu et al. (2021) as the text recognition model to compute ACC and NED.
Implementation Details. Our method is implemented using PyTorch. All experiments are conducted on a multi-GPU setup with four NVIDIA A100 GPUs, each equipped with 80GB of memory. We set the batch size to 64 and use the Adam optimizer Kingma and Ba (2014) with a learning rate of $1\times 10^{-5}$. In the loss function, the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are empirically set to 1, 1, and 0.02, respectively.
Table 1: Quantitative comparison for the synthetic dataset on CTR-TSR-Test Yu et al. (2021) with different methods.